OpenStudy (anonymous):

Evaluate the limit as h goes toward zero for the following equation:

OpenStudy (anonymous):

\[\frac{ 7\cos(\frac{ \pi }{ 6 }+h) - 7\cos(\frac{ \pi }{ 6 }) }{ h }\]

OpenStudy (anonymous):

I tried solving it out and got 0, but this isn't correct.

zepdrix (zepdrix):

So this is our limit definition for the derivative, ya? :)\[\large\rm \frac{d}{dx}7\cos(x)|_{x=\pi/6}\quad=\lim_{h\to0} \frac{ 7\cos(\frac{ \pi }{ 6 }+h) - 7\cos(\frac{ \pi }{ 6 }) }{ h }\]So if you know your derivative shortcut rules, you can avoid all the mess. Do you need to solve this through limit though? I'm trying to remember the trick for this one :)

OpenStudy (anonymous):

Yep! I need to solve it out to see if it equals one of these answers: a) -7 b) -7/2 c) 7sqrt(3)/2 d) doesn't exist

zepdrix (zepdrix):

Oh oh now I remember :) We need to apply our Angle Sum Formula:\[\large\rm \cos\left(a+b\right)=\cos(a)\cos(b)-\sin(a)\sin(b)\]Yah, that'll get us on the right track.

OpenStudy (anonymous):

@zepdrix Sorry I ditched you in the middle of the question zep. I got unfairly banned, no response so far though.

OpenStudy (anonymous):

@zepdrix Can you quickly help before this account gets banned too?

OpenStudy (anonymous):

http://openstudy.com/study#/updates/567ef3dbe4b032ed60ddcf6a It might be a bit slow though

OpenStudy (anonymous):

I can't take the derivative of cosine to get 7sin(x) and then solve from there?

zepdrix (zepdrix):

\[\large\rm \lim_{h\to0}\frac{7\color{orangered}{\cos\left(\frac{\pi}{6}+h\right)}-\cos\left(\frac{\pi}{6}\right)}{h}\]Becomes,\[\large\rm \lim_{h\to0}\frac{7\left[\color{orangered}{\cos\left(\frac{\pi}{6}\right)\cos\left(h\right)-\sin\left(\frac{\pi}{6}\right)\sin\left(h\right)}\right]-7\cos\left(\frac{\pi}{6}\right)}{h}\]Sorry it took me a sec to write all that out XD That step make sense hopefully?

zepdrix (zepdrix):

Yes, that is your other, more straight forward option :)

zepdrix (zepdrix):

Derivative of cos(x) is -sin(x) though, don't forget the minus :)

OpenStudy (anonymous):

Why cos(pi/6)cos(h)-sin(pi/6)sin(h)?

zepdrix (zepdrix):

If you're allowed to use your derivative shortcuts, then we don't need to worry about it. But that would be an important step in solving it through the limit.

zepdrix (zepdrix):

After that we would group things up in a clever way, and apply some limit identities.

zepdrix (zepdrix):

I applied the Cosine Angle-Sum Identity, too confusing? :)

zepdrix (zepdrix):

Or just confusing `why` we would need that?